# Decentralization of Ethereum’s Builder Market

### v4

@@ 1.1 Proposer Loss in the MEV-Boost Auctions
Auction data.   A practical challenge in quantifying the proposer loss in past MEV-Boost auctions is the lack of data. Only the winning bids are recorded in the blockchain, but to compute the losses, we also need losing bids. The key ingredient that enabled our study is a large-scale auction dataset we curated since 2022: we collected 8.79 billion partial bids [^1] Partial and full bids are to be defined in Sec. 3.1 of auctions from September 2022 to July 2024; in addition, in collaboration with ultra sound relay [15], we obtained full bids for a total of 301,479 auctions from April 2023 to December 2023. From the auction dataset, we can compute the true value of all bids.
@@ 1.2 Private Order Flows
Not every provider is equally important, and some providers who generate profitable order flows can significantly impact builders. Thus, we propose a metric to identify important providers and investigate their implications.
Identifying pivotal private order flow providers.   To formally capture the intuition of “important providers”, we defined pivotal level to quantify the sustained impact of a given provider. A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ were not available. In other words, pivotal providers are necessary for the winner to win. The pivotal level of a provider $P$ is the percentage of MEV-Boost auctions in which $P$ is pivotal. Other metrics, calculated based on each provider’s contribution to the builders’ blocks [18, 19], may include temporarily influential providers, whereas this metric identifies sustainedly influential providers, resulting in fewer false positives.
@@
The reputability of a builder is accessed per their specific rules. While top builders usually pass the assessments, it is harder for new builders to enter the market.
Exclusive pivotal providers.   A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders. Besides the well-known exclusive relationships (known as “integration” [20]), we also identified that three of the top five pivotal providers, previously not known to have integration, began integrating with top builders in 2024. Our finding shows that the integration of these three pivotal providers accounted for at least 45% of the total proposer losses in 2024.
@@ 1.3 Mitigation
  • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest. Our dataset of 11 billion partial bids since 2022 is the largest scale known to us.
  • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions. The findings in this paper can serve as empirical evidence for future theoretical interrogation.
@@ 2.1 MEV-Boost Auctions
  • Builders submit bids to relays throughout the slot. For this paper, we denote a bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key. Bids are signed by the builder but we omit the signature for clarity. Relays verify that $B$ indeed pays the proposer at least $BV$.
@@ 2.2 MEV Supply Chain
Builder’s income.   We assume a builder $p$’s income from building a block $B$ is $EV_{p}(B)$, or $EV(B)$ for short since the builder identity is usually encoded in the block. We compute EV following a similar approach as that in MEV analysis tools [41, 42] and previous works [43, 44]. Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees [45] and direct transfers [46] to the builder’s addresses minus the refund to other addresses. We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
@@ 3 Dataset
We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to October 2024, and a dataset of full bids from April to December 2023. The full bids are obtained from the ultra sound relay [15]. Recall from Sec. 2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc. To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. I summarizes the datasets.
@@ 3.1 Data Collection and Validation
  • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx. A for details).
@@
  • Private transaction hashes: We use data from Mempool Guru [58] to label transactions that bypassed the public mempool as private. We believe these labels have already been cross-validated since the private transactions are collected by seven distributed nodes.
  • Transaction sources: we are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through. We identified transaction sources using ground truth data (as described in Appx. A).
@@ 4.1 Competitiveness of MEV-Boost Auctions
To compute $TV(B)$ of a given block $B$, we follow the standard approach in MEV analysis tools [41, 42] and previous works [43, 44], though we built our own analysis toolchain and optimized its performance to handle the large volume of historical auction data (326.6 million full bids). We compute $EV(B)$ using a modified Ethereum execution client (Reth [47]) to execute transactions in $B$ and compute the net ETH balance changes of the builder’s addresses. Most builders’ addresses are well-known partly because they need to build up a reputation. We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec. 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
@@
Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value. $CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, thus satisfying the definition of a competitive auction, while $CI(s)<0$ means there is insufficient competition.
@@
Analysis of uncompetitiveness and inefficiency.   Fig. 4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the our dataset. Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
@@
Loss from uncompetiveness.   For an uncompetitive auction, the loss from uncompetitiveness is the difference between the second-highest true value and the actual winning bid value, representing the loss a validator incurs compared to what it would have received in a competitive auction. We computed this loss using our dataset, and the results are shown in Tab. II. The loss from uncompetitiveness ranges from 0.5% to 1.7% during our study periods.
@@
Time Slots Profits (ETH) Losses-un (ETH) (%) Losses-in (ETH) (%)
@@ 4.2 Inequality in Block-building Capacities
The previous section shows that the MEV-Boost auctions are competitive in most cases. However, a proposer may still incur loss due to a significant block-building capacity gap between builders. In MEV Boost auctions, this quantity is captured by the true values of the builder’s bids.
@@
Figure 6: Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders. The QCD measures the inequality of block-build capacity across builders. This figure reveals that such inequality worsens as MEV increases. Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
@@
Loss from inequality.   For each auction, the loss from inequality is the difference between the highest and second-highest true values, reflecting the potential loss a validator may face due to unequal block-building capacities, as a competitive auction only ensures the winning bid slightly exceeds the second-highest true value.
As shown in Tab. II, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset. Comparing the two types of loss, we find that inequality is the primary source of proposers’ loss.
@@ 5.1 Pivotal Private Order Flow Providers
Accessibility by multiple builders.   Among the top five pivotal providers, MEV-Share and MEV Blocker are public-facing services with clear documentation and open communication channels. This enables us to analyze their statements regarding how their order flows are distributed to different builders—both claim that every connected builder receives the order flows [35, 36]. Although their order flows are not accessible to all builders (as summarized in Tab. V), top builders such as Beaver, Titan, and Rsync have access to them. Ideally, competition among these top builders will ensure that the value from these order flows is ultimately captured by the proposers.
Notably, while these order flows are distributed equally to all connected builders, inequality still exists. Order flow providers may exclude less reputable builders due to trust concerns. Although this exclusion may not harm proposers, it could exacerbate builder centralization, as new builders must expend significant effort to meet the requirements imposed by these pivotal providers. Due to page constraints, we provide a more detailed discussion of the trust barrier and how it prevents builder decentralization in Appx. E.2.
@@ 5.2 Exclusive Pivotal Providers
before 2024 2024
SCP CEX-DEX arbitrage Beaver 18,509 4,149.3  (21.0) Banana Gun Sniper Titan 33,748 19,401.3  (37.1)
Searcher: 0x57c1… CEX-DEX arbitrage Addr: 0x4737… 7,166 1,037.3  (5.2) SCP CEX-DEX arbitrage Beaver 10,087 3,249.8  (6.2)
Wintermute CEX-DEX arbitrage Rsync 6,555 553.1  (2.8) jaredfromsubway.eth Sandwich Beaver 20,213 2,575.4  (5.0)
Searcher: 0xf8b7… CEX-DEX arbitrage Manta 4,606 428.2  (2.2) Searcher: 0x6980… Atomic arbitrage I can haz block? 513 2,484.9  (4.8)
Searcher: 0x9dd8… Atomic arbitrage Addr: 0x3bee… 1,600 411.5  (2.6) Searcher: 0x6454… Liquidation Addr: 0x3bee… 793 1,229.6  (2.4)
Searcher: 0x761d… Atomic arbitrage builder0x69 12 287.0  (1.5) c0ffeebabe.eth Liquidation Ty For The Block 367 1,003.6  (1.9)
Searcher: 0x6454… Liquidation Addr: 0x3bee… 539 262.5  (1.5) Banana Gun Sniper Beaver 2,430 987.4  (1.9)
Searcher: 0x5e51… Atomic arbitrage eth-builder 266 223.2  (1.1) WINTERMUTE CEX-DEX arbitrage Rsync 6,503 976.8  (1.9)
Searcher: 0xeeaa… Atomic arbitrage eth-builder 51 194.1  (1.0) Maestro Sniper Beaver 10,103 520.3  (1.0)
Searcher: 0xe20c… Atomic arbitrage Addr: 0x5638… 579 168.5  (0.9) Searcher: 0x77d5… Atomic arbitrage Rsync 782 345.3  (0.7)
@@
Compared to the exclusive order flows before 2024, a more concerning problem is the formation of more exclusive relationships between top builders and top providers. This also confirms our conjecture. For instance, Banana Gun exclusively sends order flows to Titan in approximately 33K blocks, accounting for 37.1% of the estimated proposer losses covered in our study. Additionally, two other top-5 pivotal providers, jaredfromsubway.eth and Maestro, also exclusively send order flows to Beaver, contributing to 6% of the losses. In contrast, the total losses caused by the integrated CEX-DEX arbitrage searchers remain relatively similar before and during 2024. Notably, searcher 0x57c1… and searcher 0xf8b7… stopped operating in 2023. These results effectively explain that compared to the previous known integration, the emergence of new integration accounts for the majority of the proposer losses in 2024.
@@ 6.3 Limitations of Our Approach and Data
We compute the true values of bids following a standard approach as [41, 42, 43, 44]. This approach relies on identifying builder-controlled addresses, and missing addresses could cause inaccuracies. We cross-validate with multiple sources to ensure collected addresses are the most comprehensive to date (see Sec. 3.1). It may underestimate values by ignoring off-chain payments, or overestimate profits if builders refund providers off-chain. A builder with large off-chain payments could bid higher than the computed true value, appearing as an extremely large subsidy. However, we observe that there are no excessively large subsidies, and most builders bid close to our computed true values in our study of competitiveness and inequality, suggesting that off-chain payments are not significant. The potential refunds to providers do not affect the computation of proposers’ losses because this portion is not allocated to proposers, regardless of the volume of refunds. Due to page constraints, we discuss the potential implications of the limitations in Appx. F.

### v5

@@ 1.1 Proposer Loss in the MEV-Boost Auctions
Auction data.   A practical challenge in quantifying the proposer loss in past MEV-Boost auctions is the lack of data. Only the winning bids are recorded in the blockchain, but to compute the losses, we also need losing bids. The key ingredient that enabled our study is a large-scale auction dataset we curated since 2022: we collected approximately 12 billion partial bids [^1] Partial and full bids are to be defined in Sec. 3.1 of auctions from September 2022 to October 2024; in addition, in collaboration with ultra sound relay [15], we obtained full bids for a total of 301,479 auctions from April 2023 to December 2023. From the auction dataset, we can compute the true value of all bids.
@@ 1.2 Private Order Flows
Not every provider is equally important, and some providers that generate profitable order flows can significantly impact builders. Thus, we propose a metric to identify important providers and investigate their implications.
Identifying pivotal private order flow providers.   To formally capture the intuition of “important providers”, we defined pivotal level to quantify the sustained impact of a given provider. A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ not been available. In other words, pivotal providers are necessary for the winner to win. The pivotal level of a provider $P$ is the percentage of MEV-Boost auctions in which $P$ is pivotal. Other metrics, calculated based on each provider’s contribution to the builders’ blocks [18, 19], may include temporarily influential providers, whereas this metric identifies sustainedly influential providers, resulting in fewer false positives.
@@
The reputability of a builder is assessed per their specific rules. While top builders usually pass the assessments, it is harder for new builders to enter the market.
Exclusive pivotal providers.   A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders. Besides the well-known exclusive relationships (known as “integration” [20]), we also identified that three of the top five pivotal providers, previously not known to have integration, began integrating with top builders in 2024. Our findings show that the integration of these three pivotal providers accounted for significant proposer losses in 2024.
@@ 1.3 Mitigation
  • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest. Our dataset of approximately 12 billion partial bids since 2022 is the largest scale known to us. We have open-sourced the code and datasets at [24] and [25].
  • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions.
    The findings in this paper can serve as empirical evidence for future theoretical interrogation.
@@ 2.1 MEV-Boost Auctions
  • Builders submit bids to relays throughout the slot. For this paper, we denote a full bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key. Bids are signed by the builder, but we omit the signature for clarity. Relays verify that $B$ indeed pays the proposer at least $BV$.
@@ 2.2 MEV Supply Chain
Builder’s income.   We assume a builder $p$’s income from building a block $B$ is $EV_{p}(B)$, or $EV(B)$ for short since the builder identity is usually encoded in the block. We compute EV following a similar approach to that in MEV analysis tools [44, 45] and previous works [46, 47]. Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees [48] and direct transfers [49] to the builder’s addresses, minus the payment to the proposer and refund to other addresses. We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
@@ 3 Dataset
We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to October 2024, and a dataset of full bids from April to December 2023. The full bids are obtained from the ultra sound relay [15]. Recall from Sec. 2.1 that a bid consists of a block $B$ and a bid value $BV$, while a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$, where $metadata$ includes parent hash, gas limit, gas used, etc. To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. I summarizes the datasets.
@@ 3.1 Data Collection and Validation
  • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx. B for details).
@@
  • Private transaction hashes: We use data from Mempool Guru [61] to label transactions that bypassed the public mempool as private. We believe these labels have already been cross-validated, as the private transactions were not collected by any of the seven distributed nodes.
  • Transaction sources: We are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through. We identified transaction sources using ground truth data (as described in Appx. B).
@@ 4.1 Competitiveness of MEV-Boost Auctions
Thanks to the full bids dataset, we can compute not only the transactions that appear in on-chain blocks but also those included in other bids submitted in the auction. To compute $TV(B)$ of a given block $B$, we follow the standard approach in MEV analysis tools [44, 45] and previous works [46, 47], though we built our own analysis toolchain and optimized its performance to handle the large volume of historical auction data (326.6 million full bids). We compute $EV(B)$ using a modified Ethereum execution client (Reth [50]) to execute transactions in $B$ and compute the net ETH balance changes of the builder’s addresses. Most builders’ addresses are well-known partly because they need to build up a reputation. We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec. 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
@@
Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value. We define an auction as competitive if $CI(s)\geq 0$, meaning the winning bid is at least as high as the second-highest true value; if $CI(s)<0$, the auction is uncompetitive.
@@
Analysis of uncompetitiveness and inefficiency.   Fig. 4 shows the calculated CI and EI distribution across all MEV-Boost auctions in our dataset. Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
@@
Loss from uncompetiveness.   For an uncompetitive MEV-Boost auction in slot $s$, let $TV(p_{2},s)$ denote the second-highest true value and $BV_{w}(s)$ denote the winning bid. The loss from uncompetitiveness is defined as
$L_{\text{uncomp.}}(s)=TV(p_{2},s)-BV_{w}(s),$
representing the loss a proposer incurs compared to what it would have received in a competitive auction. We computed this loss using our dataset, and the results are shown in Tab. II. The loss from uncompetitiveness ranges from 0.5% to 1.7% during our study periods.
@@
Time Slots Profits (ETH) $L_{\text{uncomp.}}$ (ETH) (%) $L_{\text{inequal.}}$ (ETH) (%)
@@ 4.2 Inequality in Block-building Capacities
The previous section shows that the MEV-Boost auctions are competitive in most cases. However, a proposer may still incur loss due to a significant block-building capacity gap between builders. In MEV-Boost auctions, this quantity is captured by the true values of the builder’s bids.
@@
Figure 6: Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders. The QCD measures the inequality of block-building capacity across builders. This figure reveals that such inequality worsens as MEV increases. Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
@@
Loss from inequality.   For the auction in slot $s$, let $TV(p_{1},s)$ and $TV(p_{2},s)$ denote the highest and second-highest true values, respectively. The loss from inequality is defined as
$L_{\text{inequal.}}(s)=TV(p_{1},s)-TV(p_{2},s),$
reflecting the potential loss a proposer may face due to unequal block-building capacities, since a competitive auction only ensures the winning bid slightly exceeds the second-highest true value. As shown in Tab. II, the loss from inequality ranges from 5.6% to 11.5% over the periods in our dataset. Comparing the two types of loss, we find that inequality is the primary source of proposers’ loss.
@@ 5.1 Pivotal Private Order Flow Providers
Accessibility by multiple builders.   Among the top five pivotal providers, MEV-Share and MEV Blocker are public-facing services with clear documentation and open communication channels. This enables us to analyze their statements regarding how their order flows are distributed to different builders—both claim that every connected builder receives the order flows [38, 39]. Although the order flows are not accessible to all builders (as summarized in Tab. V, presented in Sec. E.2), top builders such as Beaver, Titan, and Rsync have access to them. Ideally, competition among these top builders will ensure that the value from these order flows is ultimately captured by the proposers.
Notably, while these order flows are distributed equally to all connected builders, inequality still exists. Order flow providers may exclude less reputable builders due to trust concerns because builders are at an advantage and malicious builders can harm providers by, e.g., sandwich attacks [62] and imitation attacks [63]. Although this exclusion may not harm proposers, it could exacerbate builder centralization, as new builders must expend significant effort to meet the requirements imposed by these pivotal providers. Due to page constraints, we provide a more detailed discussion of the trust barrier and how it prevents builder decentralization in Appx. E.2.
@@ 5.2 Exclusive Pivotal Providers
before 2024 2024
SCP CEX-DEX arbitrage Beaver 42,639 4,566.0  (22.7) Banana Gun Sniper Titan 34,442 17,488.1  (32.6)
Searcher: 0x57c1… CEX-DEX arbitrage Addr: 0x4737… 12,220 1,088.7  (5.4) SCP CEX-DEX arbitrage Beaver 13,148 3,077.8  (5.7)
Wintermute CEX-DEX arbitrage Rsync 13,823 642.6  (3.2) Searcher: 0x6980… Atomic arbitrage I can haz block? 516 2,495.7  (4.7)
Searcher: 0xf8b7… CEX-DEX arbitrage Manta 9,699 473.9  (2.4) jaredfromsubway.eth Sandwich Beaver 22,230 2,480.5  (4.6)
Searcher: 0x9dd8… Atomic arbitrage Addr: 0x3bee… 1,826 424.9  (2.1) Searcher: 0x6454… Liquidation Addr: 0x3bee… 809 1,231.8  (2.3)
Searcher: 0x761d… Atomic arbitrage builder0x69 19 287.5  (1.4) WINTERMUTE CEX-DEX arbitrage Rsync 8,586 1,038.3  (1.9)
Searcher: 0x6454… Liquidation Addr: 0x3bee… 456 246.6  (1.2) c0ffeebabe.eth Liquidation Ty For The Block 368 1,003.6  (1.9)
Searcher: 0x5e51… Atomic arbitrage eth-builder 370 226.5  (1.1) Banana Gun Sniper Beaver 2,424 920.5  (1.7)
Searcher: 0xeeaa… Atomic arbitrage eth-builder 67 195.6  (1.0) Maestro Sniper Beaver 10,587 439.7  (0.8)
Searcher: 0xe20c… Atomic arbitrage Addr: 0x5638… 592 168.9  (0.8) Searcher: 0x77d5… Atomic arbitrage Rsync 797 333.3  (0.6)
@@
Compared to the exclusive order flows before 2024, a more concerning problem is the formation of more exclusive relationships between top builders and top providers. This also confirms our conjecture. For instance, Banana Gun exclusively sends order flows to Titan in approximately 34K blocks, accounting for 32.6% of the estimated proposer losses covered in our study. Additionally, two other top-5 pivotal providers, jaredfromsubway.eth and Maestro, also exclusively send order flows to Beaver, contributing to 5.4% of the losses. In contrast, the total losses caused by the integrated CEX-DEX arbitrage searchers remain relatively similar before and during 2024. Notably, searcher 0x57c1… and searcher 0xf8b7… stopped operating in 2023. These results effectively explain that compared to the previous known integration, the emergence of new integration accounts for the majority of the proposer losses in 2024.
@@ 6.3 Limitations of Our Approach and Data
We compute the true values of bids following a standard approach as [44, 45, 46, 47]. This approach relies on identifying builder-controlled addresses, and missing addresses could cause inaccuracies. We cross-validate with multiple sources to ensure collected addresses are the most comprehensive to date (see Sec. 3.1). It may underestimate values by ignoring off-chain payments, or overestimate profits if builders refund providers off-chain. A builder with large off-chain payments could bid higher than the computed true value, appearing as an extremely large subsidy. This could lead to the misclassification of an efficient auction as inefficient due to subsidization. However, we observe that there are no excessively large subsidies, and most builders bid close to our computed true values in our study of competitiveness and inequality, suggesting that off-chain payments are not significant. While underestimating the winner’s true value does not affect our competitiveness assessment, underestimating the second-highest true value could cause an uncompetitive auction to be misclassified as competitive. Nevertheless, this scenario is rare. In practice, builders with higher true values would have an incentive to bid higher to win, making such misclassifications unlikely.
The potential refunds to providers do not affect the computation of proposers’ losses because this portion is not allocated to proposers, regardless of the volume of refunds. However, overestimating the second-highest true value could lead us to misclassify a competitive auction as uncompetitive. Our study shows that large-scale integration between top builders and providers—where such refunds are more likely—began in 2024, after the period covered by our evaluation of competitiveness and efficiency.
We might also miss pivotal providers due to incomplete coverage of transaction sources. For example, BloXroute (a widely used channel) does not publish the list of transactions they processed. As a result, some providers may be missing from our analysis. However, there will be no false positives because each provider attribution is either directly from official sources or through known searcher addresses (see Sec. 3). This limitation leads to an underestimation of the significance of providers, but not an overestimation. Similarly, because partial bids are self-reported by relays, underestimating the second-highest bid could cause us to underestimate a provider’s pivotal level.